# SPA-Bench: A Comprehensive Benchmark for SmartPhone Agent Evaluation

### v1

@@ 5.2 Success Detection
Single-App Success Detection. We employ a coarse-to-fine success detection pipeline that uses key component matching followed by MLLM evaluation. As shown in Figure 4, for each agent-task pair, the pipeline first applies coarse detection, filtering out trajectories irrelevant to the task. If passed, fine detection follows, using an MLLM evaluator for final success determination. We compared our single-app success detection approach with human evaluations and found it achieves an F1 score of 0.926 for English tasks and 0.884 for Chinese tasks. Further details on the single-app success detection and its performance can be found in Appendix D.
@@ 6 Experiments
Since some agents have multiple variants, we provided detailed agent configurations in Appendix F.1.
@@ 6.1 Overview of Success Rate
In this paper, the success rate results were generated using the automated success detection methods described in Section 5.2, with GPT-4o serving as the MLLM.
Table 2: Success rates across all tasks and agents in this benchmark, categorised by task type. The first seven agents are based on GPT-4o, while the last four use fine-tuned MLLMs.
@@
English Chinese English Chinese
Off-the-Shelf Model (GPT-4o)
AppAgent 0.340 0.247 0 0
AutoDroid 0.327 0.187 0 0
MobileAgent 0.387 0.240 0.050 0.100
MobileAgentV2 0.433 0.440 0.100 0.100
M3A 0.640 0.447 0.200 0.100
T3A 0.487 0.380 0.100 0.100
SeeAct 0.393 0.327 0.100 0.050
Fine-tuned Model
Auto-UI 0.013 0.007 0 0
CogAgent 0.027 0.027 0 0
DigiRL 0.020 0 0 0
OdysseyAgent 0.053 0.020 0 0
@@
Comparison in Single-App Tasks. Table 2 shows the overall success rates. For single-app English tasks, M3A, T3A, and MobileAgentV2 performed the best, with success rates ranging from 0.640 to 0.433. These agents are equipped with reflection modules that help prevent them from stalling. AppAgent and AutoDroid performed less well, though they would likely had performed better with access to external knowledge documents, as in their original implementations. For single-app Chinese tasks, MobileAgentV2 outperformed T3A, while its performance was more comparable to M3A. A potential reason is that the accessibility tree layout documents used in AndroidWorld are overly complex for GPT-4o to process, compared to the OCR information used by MobileAgentV2 in Chinese apps. Generally, a decrease in success rates for Chinese tasks was observed due to the limited capabilities of (M)LLMs in Chinese. This decline was further influenced by the increased complexity of Chinese apps, which often feature more intricate layouts, frequent animations, and distracting elements like ads and pop-ups.
@@
Impact of Core Models and Input Modalities. There was a significant gap in success rates between agents using proprietary models like GPT-4o and those based on fine-tuned models. Agents based on GPT-4o consistently outperformed fine-tuned MLLMs, the latter of which often struggled to complete any tasks. This contrasts with the high action matching scores reported in prior studies Zhan & Zhang (2023); Hong et al. (2024); Bai et al. (2024); Lu et al. (2024), suggesting that fine-tuning typically focuses on generating individual actions or textual plans without adequately grounding them in real-world contexts. For instance, while certain actions (like dual-point taps) allow for flexibility in offline environments (e.g., within 14% of the screen distance from a golden standard Rawles et al. (2024b)), such approximations may lead to inaccuracies during on-device execution. On the other hand, agents using proprietary models are typically equipped with input from visual modules, such as mark-up documents and set-of-marks Yang et al. (2023a). These layout documents are sometimes incomplete, failing to capture all available UI elements on the interface. In other cases, they are unnecessarily complex for models to handle, as seen in the case of T3A mentioned above. This highlights a critical gap in grounding capabilities, which are essential for end-to-end task completion but remain challenging especially for fine-tuned models Zheng et al. (2024).
@@
Complexity and Memory Retention in Cross-App Task. For cross-app tasks, most agents, except M3A, completed no more than 4 tasks in total across both English and Chinese apps. Although M3A performed better, completing 6 out of 40 tasks, overall performance was still low, reflecting the complexity of cross-app tasks. These tasks require more steps, reasoning, and the ability to switch between apps while retaining memory of previous actions. In some cases, agents might nearly complete the task but fail in the end due to minor mistakes or missed requirements, especially in long sequences or multi-context scenarios. Even OdysseyAgent Lu et al. (2024), specifically designed for cross-app tasks, faced difficulty completing them end-to-end. It sometimes handled subtasks within a single app well but faltered when transitioning between apps, illustrating the challenge of maintaining context and reasoning across environments. These findings suggest that current agents, including the best-performing ones, struggle with multi-step cross-app tasks, often losing context or forgetting prior actions. This highlights the need for better memory mechanisms, enhanced inter-app reasoning, and advanced handling of complex, multi-context environments Shinn et al. (2023); Li et al. (2023b); Pan et al. (2024). These capabilities are essential for tasks users expect autonomous agents to manage.
@@ 6.2 Completion- and Consumption-related Metrics
Agent Success (%) Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC (%) MSR (%) Error (%) Premature (%) Overdue (%)
Off-the-Shelf Model (GPT-4o)
@@
Step Efficiency and Success Rate. As discussed in Section 6.1, agents based on off-the-shelf (M)LLMs significantly outperformed fine-tuned MLLMs. Higher success rates correlate with lower mean step ratios, indicating more efficient task completion with fewer unnecessary actions or errors. Conversely, agents facing difficult tasks tend to make more errors, even if they ultimately succeed. M3A exhibited a notably low mean step ratio of 0.92, indicating it used fewer steps than a human. This efficiency is partly due to combining actions, such as pressing “enter” immediately after typing in the search box, saving one step. Some agents also exploited shortcuts, like selecting recommended search options instead of typing keywords and pressing the search button.
@@
Enhancing Robustness through Error Handling Mechanisms. Error-handling mechanisms are crucial for improving success rates and ensuring reliable performance. Agents lacking these mechanisms were more vulnerable to failure or premature termination when execution errors occur. Common issues include parsing errors due to invalid formats or difficulties translating model outputs into valid actions. Some agents failed when necessary inputs (e.g., XML files), could not be accessed. These failures highlight the need for better error detection and recovery strategies, allowing agents to correct mistakes and improve their overall success rates.
Limitations in Cost and Efficiency for Real-World Use. While fine-tuned MLLM-based agents do not incur token costs and their execution time varies with device power, their low success rates make them impractical for deployment. Among agents based on GPT-4o, AutoDroid is the most cost-effective, using only \$0.008 per step due to its text-based input. However, it has a long execution time (34 seconds per step) and a success rate of only 0.327. M3A and T3A, though faster (under 20 seconds per step) and more successful, have higher token costs at around \$0.10 per step due to the complexity of inputs generated by UI elements. MobileAgentV2, while more affordable at \$0.067 per step, suffers from a complex visual perception pipeline, resulting in the longest execution time (56.1 seconds per step). These results highlight the trade-off between efficiency and effectiveness. Agents like T3A, despite achieving relatively high success rates and faster execution times, still fall short of human-level usability due to their monetary cost. Such limitations stem from two major factors. One is the delay between UI information collection and action execution, which can cause inaccuracies especially when dynamic content appears. The other is the agents’ slower speeds and higher costs compared to human users. Users are unlikely to rely on an autonomous agent to complete a task if they have to wait for extended periods or pay several dollars, especially when they could complete it in a few steps themselves.
Performance Variation Difficulty Levels and Open-Source Replacements. We also compared performance across tasks at each difficulty level and replace GPT-4o with open-source (M)LLMs. As expected, easier tasks were executed more successfully, and agents using open-source models showed a performance gap compared to their GPT-4o counterparts. More details can be found in Appendix F.3 and Appendix F.4.
Case Study. In Appendix G, three case studies are presented to illustrate representative scenarios of task execution by agents. These include: (1) an invalid action taken by AppAgent due to misinterpretation of the UI structure in the XML file, (2) a dynamically changing screen without any action execution, repetitive actions due to the lack of reflection, and unrelated behaviours to the task description in MobileAgent, and (3) the shortcut strategy employed by M3A.
@@ 6.3 Key Insights
First, integrating more advanced visual perception modules is essential for enhancing agents’ understanding of complex UI layouts and precise action grounding across various scenarios. Although agents using accessibility trees and OCR have shown relatively good performance in English tasks, their effectiveness is still limited in Chinese tasks, which often feature more visually complex and dynamic content. Currently, some agents struggle to ground actions in these dynamic environments, often failing to recognise actionable elements or map generated actions to the correct coordinates. Future designs should focus on building more robust visual models that can accurately interpret these environments and perform end-to-end task completion in interactive settings.
@@
Finally, execution time and cost need to be optimised for real-world deployment. Agents such as MobileAgentV2, which rely on multiple modules, need to reduce overhead and streamline execution to minimise task completion time. MLLM-based agents, in contrast to T3A, may also focus on reducing input context size to lower token costs while preserving critical information for task completion. A hybrid model approach that combines the speed and efficiency of lightweight models with the robustness of more complex ones could provide a promising solution for balancing performance and cost in real-world applications.
@@ 7 Conclusion
In this paper, we introduced SPA-Bench, a comprehensive benchmark for evaluating smartphone agents across diverse tasks. The evaluation covers English and Chinese apps, single-app and cross-app scenarios, and varying difficulty levels. Our experiments reveal that even the best-performing agents can complete less than 70% of tasks successfully, and there are significant performance gaps between agents using proprietary (M)LLMs and those relying on open-source or fine-tuned models, particularly in action grounding and generalisation within complex Chinese apps. While some agents excel in simpler tasks, their long execution times and high costs limit their practicality for real-world use. Our findings highlight the need for better memory mechanisms, robust error handling, accurate self-evaluator, improved integration of reasoning with UI understanding, and optimising execution time and cost for real-world deployment. Additionally, agents based on fine-tuned models should be adapted to diverse scenarios and focus on long-sequence decision-making rather than isolated actions. By developing SPA-Bench as a fair and scalable benchmark, we aim to accelerate the development of more efficient, practical, and user-friendly smartphone agents.
@@ Appendix A Limitation and Future Work
Given that constructing tasks is both time-consuming and resource-intensive, SPA-Bench currently includes 300 single-app tasks and 40 cross-app tasks, evenly split between English and Chinese. We plan to expand the scope of our task collection and increase the diversity in task presentation (e.g., by adding vague task descriptions and mimicking different human tones). Since some apps are difficult to operate using emulators, we also aim to design tasks that can be more easily experimented with. Additionally, we will execute experiments multiple times to ensure robustness.
@@ B.2.2 Single-App Chinese Tasks
{CJK}UTF8gbsn哔哩哔哩 3 9 {CJK}UTF8gbsn评论, 哈哈 {CJK}UTF8gbsn浏览个人消息通知,挑选一个聊天,查看好友动态,点赞好友的一条动态,编辑评论“哈哈”。
@@
{CJK}UTF8gbsn小红书 3 11 {CJK}UTF8gbsn发送, 很有用, 谢谢 {CJK}UTF8gbsn搜索一种名为“减肥食谱”的笔记,按照热度排序,观看其中一个热度最高的笔记,点赞该笔记,收藏该笔记,然后编辑评论“很有用,谢谢“,停留在评论发送页面不要发送。
@@ B.2.4 Cross-App Chinese Tasks
Multi Apps {CJK}UTF8gbsn华为浏览器, bilibili, QQ 2 14 {CJK}UTF8gbsn在华为浏览器中搜索“贝塞斯达最成功的游戏是什么”,然后在bilibili搜索任意一个有关该游戏的视频,观看视频并分享到QQ空间
@@
Social Sharing {CJK}UTF8gbsnbilibili, QQ 1 9 {CJK}UTF8gbsn在bilibili中搜索“自制关卡 胆小菇之梦”,点击进入任意一个视频,分享该视频到qq空间

### v2

@@ 5.2 Success Detection
Single-App Success Detection. We employ a coarse-to-fine success detection pipeline that uses key component matching followed by MLLM evaluation. As shown in Figure 4, for each agent-task pair, the pipeline first applies coarse detection using the annotated key components introduced in Section 3.2, filtering out trajectories irrelevant to the task. If passed, fine detection follows, using an MLLM evaluator for final success determination. This approach is motivated by the need to balance scalability and cost efficiency, addressing the limitations of relying on extensive human labour or expensive MLLM-based evaluations in large-scale benchmarks. We compared our single-app success detection approach with human evaluations and found it achieves an F1 score of 0.926 for English tasks and 0.884 for Chinese tasks. Further details on the single-app success detection and its performance can be found in Appendix D.
@@ 6 Experiments
In this paper, the success rate results were derived using the automated success detection methods outlined in Section 5.2, with GPT-4o serving as the MLLM. To account for agents with multiple variants, detailed configurations for each agent are provided in Appendix F.1. Furthermore, case studies illustrating various agent behaviors are presented in Appendix H.
@@ 6.1 Overview of Success Rate
Table 2 shows the overall success rates. Notably, M3A consistently achieved the highest success rates across all task sets. We found that agents generally performed English tasks better than Chinese tasks, agents with the agentic workflow outperformed those categorised as agent-as-a-model, and cross-app tasks were more challenging than single-app tasks for agents.
Table 2: Success rates across all tasks and agents in this benchmark, categorised by task type. The first seven agents fall under the category of agentic workflow, while the last four belong to agent-as-a-model. AutoDroid was tested only on single-app tasks as its agent framework, Droidbot (Li et al., 2017), supports only these tasks.
@@
Overall English Chinese Overall English Chinese
Agentic Workflow (GPT-4o)
AppAgent 0.294 0.340 0.247 0 0 0
AutoDroid 0.257 0.327 0.187 - - -
MobileAgent 0.314 0.387 0.240 0.075 0.050 0.100
MobileAgentV2 0.437 0.433 0.440 0.100 0.100 0.100
M3A 0.544 0.640 0.447 0.150 0.200 0.100
T3A 0.434 0.487 0.380 0.100 0.100 0.100
SeeAct 0.360 0.393 0.327 0.075 0.100 0.050
Agent-as-a-Model
Auto-UI 0.010 0.013 0.007 0 0 0
CogAgent 0.027 0.027 0.027 0 0 0
DigiRL 0.010 0.020 0 0 0 0
OdysseyAgent 0.037 0.053 0.020 0 0 0
@@
Comparison in Single-App Tasks. For single-app English tasks, M3A, T3A, and MobileAgentV2 were the best-performing ones, with success rates ranging from 0.640 to 0.433. These agents are equipped with reflection modules that help prevent them from stalling. AppAgent and AutoDroid performed less well, though they would likely had performed better with access to external knowledge documents, as in their original implementations. For single-app Chinese tasks, MobileAgentV2 outperformed T3A, while its performance was more comparable to M3A. A potential factor could be the overly complex accessibility (a11y) tree layout used by T3A. MobileAgentV2, relying on OCR and raw screenshots, averaged 12,400 prompt tokens per step in Chinese single-app tasks, compared to T3A’s 22,000 tokens using only a11y trees, indicating larger or more intricate structures in Chinese apps, potentially contributing to the agent’s degraded performance. A similar trend was observed in English single-app tasks, with lower token usage across both agents: 11,200 for MobileAgentV2 and 19,700 for T3A. In general, a decrease in success rates for Chinese tasks was observed due to the limited capabilities of (M)LLMs in Chinese, compounded by the increased complexity of Chinese apps. These apps often feature more intricate layouts, frequent animations, and distracting elements such as ads and pop-ups.
@@
Impact of Core Models and Input Modalities. There was a significant gap in success rates between agents using proprietary models like GPT-4o and those based on fine-tuned models. Agents following the agentic workflow significantly outperformed those in the agent-as-a-model category, the latter of which often struggled to complete any tasks. This contrasts with the high action matching scores reported in prior studies (Zhan & Zhang, 2023; Hong et al., 2024; Bai et al., 2024; Lu et al., 2024), indicating that fine-tuned agents are often optimised for generating textual actions based on fixed UI scenarios. While such optimisations may achieve high accuracy in offline environments, they often fail in dynamic, real-world settings. For example, a tap action is deemed successful if its coordinates fall within 14% of the screen distance to the ground truth (Rawles et al., 2024b), but this tolerance can cause inaccuracies with actionable elements in practice. Furthermore, reliance on predefined scenarios limits the agents’ ability to generalise to unseen UI contexts or to recover from detoured states caused by mistaken actions. On the other hand, agents utilising agentic workflow are typically equipped with input from visual modules, such as mark-up documents and set-of-marks (Yang et al., 2023a). These layout documents are sometimes incomplete, failing to capture all available UI elements on the interface. In other cases, they are unnecessarily complex for models to handle, as seen in the case of T3A mentioned above. This highlights a critical gap in grounding capabilities, which are essential for end-to-end task completion but remain challenging especially for fine-tuned models (Zheng et al., 2024).
@@
Complexity and Memory Retention in Cross-App Task. For cross-app tasks, most agents, except M3A, completed no more than 4 tasks in total across both English and Chinese apps. Although M3A performed better, completing 6 out of 40 tasks, overall performance was still low, reflecting the complexity of cross-app tasks. These tasks require more steps, reasoning, and the ability to switch between apps while retaining memory of previous actions. In some cases, agents might nearly complete the task but fail in the end due to minor mistakes or missed requirements, especially in long sequences or multi-context scenarios. Even OdysseyAgent (Lu et al., 2024), specifically designed for cross-app tasks, faced difficulty completing them end-to-end. It sometimes handled subtasks within a single app well but faltered when transitioning between apps, illustrating the challenge of maintaining context and reasoning across environments. These findings suggest that current agents, including the best-performing ones, struggle with multi-step cross-app tasks, often losing context or forgetting prior actions. This highlights the need for better memory mechanisms, enhanced inter-app reasoning, and advanced handling of complex, multi-context environments (Shinn et al., 2023; Li et al., 2023b; Pan et al., 2024). These capabilities are essential for tasks users expect autonomous agents to manage.
@@ 6.2 Completion- and Consumption-related Metrics
Agent Success Rate Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC Rate MSR Rate Error Rate Premature Rate Overdue Rate
Agentic Workflow (GPT-4o)
@@
Step Efficiency and Success Rate. As discussed in Section 6.1, agents with the agentic workflow substantially outperformed those belong to agent-as-a-model. Higher success rates correlate with lower mean step ratios, indicating more efficient task completion with fewer unnecessary actions or errors. Conversely, agents facing difficult tasks tend to make more errors, even if they ultimately succeed. M3A exhibited a notably low mean step ratio of 0.92, indicating it used fewer steps than a human. This efficiency is partly achieved through combined actions specifically defined by the agent itself, where a single action encompasses multiple operations, such as typing in the search box and pressing “enter” in one step. Agents may also exploit strategic shortcuts, such as clicking on a recommended item instead of using the search bar. Thus, both approaches allow agents to reduce the steps needed to complete a task.
@@
Enhancing Robustness through Error Handling Mechanisms. Error-handling mechanisms are crucial for improving success rates and ensuring reliable performance. Agents lacking these mechanisms were more prone to failure, which led to early termination when execution errors occurred. Common issues include parsing errors arising from the agents’ inability to correctly interpret model outputs as valid actions. For example, the output may be missing specific phrases, such as “Thought: ”, that are required by the agent’s parsing module. Some agents also encountered failures when necessary inputs (e.g., XML files) could not be accessed. These failures highlight the need for better error detection and recovery strategies, allowing agents to correct mistakes and improve their overall success rates.
Limitations in Cost and Efficiency for Real-World Use. While agents categorised as agent-as-a-model do not incur token costs and their execution time varies with device power, their low success rates make them impractical for deployment. Among agents with the agentic workflow, AutoDroid is the most cost-effective, using only \$0.008 per step due to its text-based input. However, it has a long execution time (34 seconds per step) and a success rate of only 0.327. M3A and T3A, though faster (under 20 seconds per step) and more successful, have higher token costs at around \$0.10 per step due to the complexity of inputs generated by UI elements. MobileAgentV2, while more affordable at \$0.067 per step, suffers from a complex visual perception pipeline, resulting in the longest execution time (56.1 seconds per step). These results highlight the trade-off between efficiency and effectiveness. Agents like T3A, despite achieving relatively high success rates and faster execution times, still fall short of human-level usability due to their monetary cost. Such limitations stem from two major factors. One is the delay between UI information collection and action execution, which can cause inaccuracies especially when dynamic content appears. The other is the agents’ slower speeds and higher costs compared to human users. Users are unlikely to rely on an autonomous agent to complete a task if they have to wait for extended periods or pay several dollars, especially when they could complete it in a few steps themselves.
Performance Variation Across Difficulty Levels, Open-Ended Task Experiments, and Case Studies. We compared agent performance across difficulty levels, showing that easier tasks are executed more successfully, as demonstrated in Appendix F.3. To further explore the scalability of our success detection approaches, we conducted initial experiments on “open-ended” single-app English tasks, detailed in Appendix G. In Appendix H, we present three case studies that illustrate representative scenarios of agent task execution.
@@ 6.3 Key Insights
First, integrating more advanced visual perception modules is essential for enhancing agents’ understanding of complex UI layouts and precise action grounding across various scenarios. Although agents using a11y trees and OCR have shown relatively good performance in English tasks, their effectiveness is still limited in Chinese tasks, which often feature more visually complex and dynamic content. Currently, some agents struggle to ground actions in these dynamic environments, often failing to recognise actionable elements or map generated actions to the correct coordinates. Future designs should focus on building more robust visual models that can accurately interpret these environments and perform end-to-end task completion in interactive settings.
@@
Finally, execution time and cost need to be optimised for real-world deployment. Agents such as MobileAgentV2, which rely on multiple modules, need to reduce overhead and streamline execution to minimise task completion time. MLLM-based agents, in contrast to T3A, may also focus on reducing input context size to lower token costs while preserving critical information for task completion. A hybrid model approach that combines the speed and efficiency of lightweight models with the robustness of more complex ones could provide a promising solution for balancing performance and
@@ 7 Conclusion
In this paper, we introduced SPA-Bench, a comprehensive benchmark for evaluating smartphone agents across diverse tasks. The evaluation covers English and Chinese apps, single-app and cross-app scenarios, and varying difficulty levels. Our experiments reveal that even the best-performing agents can complete less than 70% of tasks successfully, and there are significant performance gaps between agents following the agentic workflow and those in the agent-as-a-model category, particularly in action grounding and generalisation within complex Chinese apps. While some agents excel in simpler tasks, their long execution times and high costs limit their practicality for real-world use. Our findings highlight the need for better memory mechanisms, robust error handling, accurate self-evaluator, improved integration of reasoning with UI understanding, and optimising execution time and cost for real-world deployment. Additionally, agents based on fine-tuned models should be adapted to diverse scenarios and focus on long-sequence decision-making rather than isolated actions. By developing SPA-Bench as a fair and scalable benchmark, we aim to accelerate the development of more efficient, practical, and user-friendly smartphone agents.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62422605, 92370132).
@@ Appendix A Limitation and Future Work
Given that constructing tasks is both time-consuming and resource-intensive, SPA-Bench currently includes 300 single-app tasks and 40 cross-app tasks, evenly split between English and Chinese. We plan to expand the scope of our task collection and increase the diversity of task presentation (e.g., as explored in the initial examples in Appendix G, by adding vague task descriptions and mimicking various human tones). Since some apps are difficult to operate using emulators, we also aim to design tasks that can be more easily experimented with. Additionally, we will execute experiments multiple times to ensure robustness.
Regarding our agent framework, we intend to expand the scope of our research by supporting a broader range of agents. This will include, but is not limited to, locally deployable agents, cloud-based agents, and agents capable of operating other smart devices, such as Android tablets and iOS devices.
@@ B.2.2 Single-App Chinese Tasks
{CJK}UTF8gbsn哔哩哔哩 3 9 {CJK}UTF8gbsn评论, $哈$ {CJK}UTF8gbsn浏览个人消息通知,挑选一个聊天,查看好友动态,点赞好友的一条动态,编辑评论“哈哈”。
@@
{CJK}UTF8gbsn小红书 3 11 {CJK}UTF8gbsn发送, 很有用, $谢$ {CJK}UTF8gbsn搜索一种名为“减肥食谱”的笔记,按照热度排序,观看其中一个热度最高的笔记,点赞该笔记,收藏该笔记,然后编辑评论“很有用,谢谢“,停留在评论发送页面不要发送。
@@ B.2.4 Cross-App Chinese Tasks
Multi Apps {CJK}UTF8gbsn华为浏览器, bilibili, $Q$ 2 14 {CJK}UTF8gbsn在华为浏览器中搜索“贝塞斯达最成功的游戏是什么”,然后在bilibili搜索任意一个有关该游戏的视频,观看视频并分享到QQ空间
@@
Social Sharing {CJK}UTF8gbsnbilibili, $Q$ 1 9 {CJK}UTF8gbsn在bilibili中搜索“自制关卡 胆小菇之梦”,点击进入任意一个视频,分享该视频到qq空间